Using the information given below for the fictitious country "Alpha," find the national saving rate.


   Household saving = 120 Business saving = 390

     Government purchases of goods and services = 220

     Government transfers and interest payments = 100

     Tax collections = 300     GDP = 2,500

I know that national savings is 510, I just can't get the national saving rate.
Answer #1

Public saving = tax collections – government purchases – transfers and interest payments

=300-220-100

= -20

Private saving = Household saving + business saving

= 120+390

= 510

National saving = Private saving + public saving

= 510+(-20)

= 490

National saving rate = National saving/GDP*100

= 490/2500*100

= 19.6%